A vertical face ogee 1.5m is operating at a design head of 3m. Calculate the depth of flow immediately dowmstream from the weir


1
Expert's answer
2021-12-15T19:01:01-0500

Pi = Pf

7.750×106 kg•m/s = 6.750×10kg•m/s + (250 kg)v'

(7.750×106 - 6.750×10)kg•m/s= (250 kg)v'

106 m/s = 250v'

v' = (106/250) m/s

v' = 4000 m/s